# The Price of Uncertainty for Social Consensus

### v2

@@ 3.1 Proof of Improved Lower Bound
Note that this sequence is almost identical to that of the initializer except we keep going until a $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ is formed by the last two layers. The first layer in the secondary set is the same as the last layer of the initializer, each vertex of which is connected with $\frac{1}{\varepsilon}$ frozen white vertices and $\frac{1}{\varepsilon}$ frozen red vertices. Note the following observation:
@@
It is our careful construction of the initializer to get to this point combined with Proposition 3 that is the key observation needed to improve the bound in Balcan et al. (2011). We now finish the construction. Note that the $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ produced at the end of the secondary set is a starting point from which we can perform the boosting operations as defined above. Here we add a layer of $\frac{2}{\varepsilon}+1$ vertices to form an initial $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}+1}$ and then initialize $m$ to be $\frac{2}{\varepsilon}+1$. Starting from this layer, layers will increase in size by a constant multiplicative factor of $1+\frac{\varepsilon}{2}$ until we reach $c\cdot\varepsilon n$ vertices in the last layer for some small constant $c$ to be set later. The total number of vertices used in the boosting operations is $\Theta\left(\frac{c\cdot\varepsilon n}{1-\frac{1}{1+\frac{\varepsilon}{2}}}\right)=\Theta(cn)$. Combined with the number of vertices $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$ used in the initializer as well as the secondary set, we see that for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$ and a constant $c$ independent of $\varepsilon$ and $n$ that the total number of vertices used is at most $n$, which means our construction has the right number of vertices.
@@
How many bad edges does our construction yield? Starting from the first layer in the secondary set, we can turn each successive layer red in sequence so that the last two layers form a $K_{c\cdot\varepsilon n,c\cdot\varepsilon n}$ where the left side is red and the right half is white. However, this produces only $c^{2}\cdot\varepsilon^{2}n^{2}$ bad edges, which, recalling that we started with $\frac{1}{\varepsilon}$ bad edges, matches the bound $\Omega(\varepsilon^{3}n^{2})$ on the price of uncertainty in Balcan et al. (2011). It turns out we can produce many more bad edges in our construction by doing the following. Note that the last layer in our construction is white, and all layers between the frozen layers defined above and the last layer are now red. We now turn each vertex in the layer after the frozen layers white, which is a valid uncertain best response by Proposition 3. We can now turn each successive layer white in sequence until the last three layers are white, red, and white respectively. Repeatedly switching the colors of the vertices in the layer after the frozen layers and propagating uncertain best responses, we can reach a state in which the boosting operation layers alternate between white and red. Every edge between boosting operation layers is now bad, which makes $\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{1-\left(\frac{1}{1+\frac{\varepsilon}{2}}\right)^{2}}\right)=\Theta(\varepsilon n^{2})$ bad edges at the ending state. Recalling that we began with $\frac{1}{\varepsilon}$ bad edges, we have established that the price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ in the domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
@@ 4 Upper Bound
In addition to focusing only on the steps at which each vertex plays its first uncertain best response, we will also carefully construct and analyze a monovariant that restricts the degrees of new vertices and allows us to bound the number of edges in the graph.
@@ 4.1 Proof of Improved Upper Bound
Let us number the vertices in $P$ in order of their first best responses from first to last as $p_{1},p_{2},\dots,p_{m}$ for $m=|P|$. Then, for all integers $1\leq k\leq m$, define $S_{k}=S_{0}\cup\{p_{1},p_{2},\dots,p_{k}\}$. Now that we have set up our notation, we will prove some preliminary statements regarding the behavior of the consensus game.
@@
Clearly, $m=|P|<|V|=n$, so $m=O(n)$ and the upper bound holds. We will now attack the lower bound.
By construction, $V$ consists of the two disjoint subsets $S_{m}$ and $Q$. By Proposition 6, any vertices in $Q$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge. This means that we can delete these vertices along with all of their incident edges, which will never decrease the factor by which the number of bad edges in $G$ can be increased. By construction, prior to when $p_{1}$ plays its first best response, any edges between 2 vertices in $V\setminus S_{0}$ must be good. Since $Q=V\setminus S_{m}\subset V\setminus S_{0}$, and since vertices in $Q$ never change their colors, any edges between two vertices in $Q$ must always be good. By removing all such edges, we will never decrease the factor by which the number of bad edges in $G$ can be increased.
We are now left with a new set of vertices $Q^{\prime}\subset Q$ where every edge incident to a vertex in $Q^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=P$. For integers $1\leq k\leq m$, $p_{k}$ can only play its first best response if the number of its bad neighbors is at least the number of its good neighbors multiplied by $(1+\varepsilon)^{-1}$. By Proposition 6, the number of bad edges incident to $p_{k}$ is at most $|S_{k-1}|<|S_{m}|$. Right before $p_{k}$ plays its first best response, all edges between $p_{k}$ and a vertex in $Q^{\prime}$ must be good, so
$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|Q^{\prime}|$
$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|Q^{\prime}|$
$\Theta\left(|S_{m}|\right)$ $=\Theta\left(Q^{\prime}\right).$
Thus, the number of vertices in the graph has the same asymptotic growth as that of $S_{m}$. Recall that $|S_{m}|=|S_{0}|+m$. By Proposition 5, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin with a bad neighbor is at most twice this quantity. So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well.
$\Theta(n)$ $=|S_{m}|$
$\Theta(n)$ $=|S_{0}|+m$
$\Theta(n)$ $=O\left(\varepsilon^{-2}\right)+m$
@@
$\Theta(n)$ $=o\left(n\right)+m$
$\Theta(n)$ $=m.$
@@
Suppose that $A_{0}=\{3,2,2\}$, $\varepsilon=1/4$, and we wish to play a move on $z=4$. Then, $\lfloor(1+\varepsilon)z\rfloor=5$, and the possible multisets for $A_{1}$ are:
$\{3,3,1,1,1\},\quad\{3,2,1,1,1,1\},\;\;\text{and}\;\;\{2,2,1,1,1,1,1\}.$
@@
  • Pad $A_{k}$ with an sufficiently large (at least $z(1+\varepsilon)$) number of zeroes.
@@
  • Remove all zeroes from $A_{k+1}$.
@@
Let $T_{k}$ denote the number of edges between two vertices in $S_{k}$. Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum A_{k}$ denote $(\sum A_{k})-(\sum A_{k-1})$. For every regular move that we play on $A_{k-1}$ with an integer $z\in A_{k-1}$, the sum $\sum A_{k}$ grows from $\sum A_{k-1}$ by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges between two vertices in $S_{k}$ grows from $S_{k-1}$ by $z$. This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$. We have
$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum A_{k}$
$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum A_{k}$
$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum A_{k}$
@@
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum A_{m}-\sum A_{0}.$
@@
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}.$
@@
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}$
@@
By Proposition 11, $\varepsilon^{-1}\sum A_{0}$ must be $O(\varepsilon^{-1}n|S_{0}|)$. Since $\varepsilon=\omega(n^{-1/3})$, this is dominated by the $O(\varepsilon^{2}n^{2}|S_{0}|)$ term. Also, since we must have $\Omega(\varepsilon^{-1})$ starting bad edges, $S_{0}$ is $\Omega(\varepsilon^{-1})$ as well. This means that $O(\varepsilon n^{2})$ must also be $O(\varepsilon^{2}n^{2}|S_{0}|)$. Finally, $\varepsilon=\omega(n^{-1/3})$, so $O(\varepsilon n^{2})$ will dominate $\Theta(\varepsilon^{-1}n)$. We are left with the following terms on the RHS:
@@
We will now show that $T_{0}$ cannot grow faster than $O(\varepsilon^{2}n^{2}|S_{0}|)$. By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$. By Proposition 5, we start with $O(\varepsilon^{-2})$ bad edges, so $T_{0}=O(\varepsilon^{-2})$. This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega(n^{-1/3})$. Then, $O(\varepsilon n^{2})$ is $O(\varepsilon^{2}n^{2}|S_{0}|)$ because $|S_{0}|=\Omega(\varepsilon^{-1})$. Our equation finally becomes
@@
Intuitively, playing a regular move on $\sum A_{k-1}$ with a large $z$ would result in a big increase in from $\sum A_{k-1}$ to $\sum A_{k}$. However, playing such a move might also create many small elements in $A_{k}$. Having the sum $\sum A_{k}$ spread out across many small elements will make increasing $\sum A_{k}$ in future moves more difficult. We consider $\sum A_{k}^{2}$ as a way to measure how ”spread out” the multiset $A_{k}$ is, and this particular quantity is preferable over other measures because the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ can be related to $\sum A_{k-1}$.
@@
Eq. 1 is true because the elements of $E_{m}$ must be nonnegative integers, and Eq. 2 follows from the Cauchy-Schwarz inequality. We will only consider these inequalities (and not those with $m$ and 0 swapped) because we only need to prevent too many regular moves from being played with large $z$, which will naturally cause $\sum A_{k}^{2}$ to become small as $k$ increases.
We will begin by investigating the behavior of the sum of squares. Let the current multiset be $A_{k-1}$. Then, pick some element $z\in A_{k-1}$, and define $\zeta=\lfloor z(1+\varepsilon)\rfloor$. We can play a regular move on $A_{k-1}$ with $z$, which will involve incrementing $\zeta$ the elements $a_{1},a_{2},\dots,a_{\zeta}\in A_{k-1}$ to $a_{1}+1,a_{2}+1,\dots,a_{\zeta}+1\in A_{k}$. It is easy to see that the effect of such a move on the increase from $\sum A_{k-1}$ to $\sum A_{k}$ is
@@
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+\sum_{i=1}^{\zeta}\left((a_{i}+1)^{2}+a_{i}^{2}\right)$
@@
Eq. 3 and Eq. 4 together describe a single regular move. However, we can establish bounds upon these equations that are easier to analyze. Since $a_{1},a_{2},\dots,a_{\alpha}$ are all elements of $A_{k-1}$, their sum must be bounded by $\sum A_{k-1}$. By plugging this into our expression for the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$, we get
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+2\sum_{i=1}^{\zeta}a_{i}+\zeta$
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum_{i=1}^{\zeta}a_{j}-z^{2}$
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum A_{k-1}-z^{2}.$ (5)
@@
Notice that the changes described in the right hand sides of Eq. 3, Eq. 4, and Eq. 5 are all independent of the current value of $\sum A_{k-1}^{2}$. Furthermore, the only constraint on the final sum of squares $\sum A_{m}^{2}$ is a lower bound provided by Eq. 1. This means that always taking the upper bound for $\sum A_{k}$ from Eq. 5 will never decrease $\sum A_{m}$. ∎
Now equipped with Proposition 13, we actually no longer need that $z\in A_{k-1}$. In fact, we no longer even need to consider $A_{k-1}$ as a multiset anymore. For the rest of the proof, we will simply assume that $z$ is a positive real number. The next step is to find an upper bound for the change from $\sum A_{k-1}$ to $\sum A_{k}$.
@@
$\varepsilon z^{\prime}$ $=\lfloor z(1+\varepsilon)\rfloor-z$
$\varepsilon z^{\prime}$ $\leq z(1+\varepsilon)-z$
$z^{\prime}$ $\leq z.$
By (5), smaller values of $z$ yield larger values of $\sum A_{k}^{2}$. This is exactly what we wanted. ∎
@@
$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{i}\right)-\sum_{i=1}^{m}z_{i}^{2}.$
We can rewrite the $\sum A_{j}$ by plugging in Eq. 1 and Eq. 7.
$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{0}+i+\varepsilon\sum_{j=1}^{i}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
$0$ $\leq\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\varepsilon\sum_{j=1}^{m}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}$
$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum_{i=1}^{m}z_{i}^{2}$ (8)
@@
Plugging this result into Eq. 8 yields a quadratic polynomial in terms of $\sum_{i=1}^{m}z_{i}$ in the RHS.
$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
$0$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}-2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum A_{0}^{2}$
Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}z_{j}$ must be less than the larger root of this quadratic. This larger root is obtained through a tedious application of the quadratic formula:
$\sum_{j=1}^{m}z_{j}\leq\frac{2\varepsilon m+\sqrt{4\varepsilon^{2}m^{2}+4m^{-1}\sum A_{0}^{2}}}{2m^{-1}}$ (9)
@@
Suppose that $\varepsilon=\Omega(n^{-1/4})$. Then, $\sum_{i=1}^{m}z_{j}$ is $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
@@
If we substitute asymptotic behavior into the RHS of (9), we get

### v3

@@ 3.1 Proof of Improved Lower Bound
Note that this sequence is almost identical to that of the initializer, except we keep going until a $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ is formed by the last two layers. The first layer in the secondary set is the same as the last layer of the initializer, each vertex of which is connected with $\frac{1}{\varepsilon}$ frozen blue vertices and $\frac{1}{\varepsilon}$ frozen red vertices. We now have the following observation:
@@
Our careful construction of the initializer and Proposition 3 are the key observations needed to improve the bound in Balcan et al. (2011). We will now finish the construction. Note that the $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ produced at the end of the secondary set is a starting point from which we can perform the boosting operations as defined above. Here we add a layer of $\frac{2}{\varepsilon}+1$ vertices to form an initial $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}+1}$, and then initialize $m$ to be $\frac{2}{\varepsilon}+1$. Starting from this layer, additional layers will increase in size by a constant multiplicative factor of $1+\frac{\varepsilon}{2}$ until we reach $c\cdot\varepsilon n$ vertices in the last layer for some small constant $c$ to be set later. The total number of vertices used in the boosting operations is
$\Theta\left(\frac{c\cdot\varepsilon n}{1-\frac{1}{1+\frac{\varepsilon}{2}}}\right)=\Theta\left(\frac{c\cdot\varepsilon n}{\frac{\varepsilon}{2+\varepsilon}}\right)=\Theta(cn).$
Combined with the number of vertices $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$ used in the initializer as well as the secondary set, we see that for $\varepsilon=\tilde{\Omega}(n^{-1/2})$ and a constant $c$ independent of $\varepsilon$ and $n$ that the total number of vertices used is at most $n$, which means our construction has the right number of vertices.
@@
How many bad edges does our construction yield? Starting from the first layer in the secondary set, we can turn each successive layer red in sequence so that the last two layers form a $K_{c\cdot\varepsilon n,c\cdot\varepsilon n}$ where the left side is red and the right half is blue. However, this produces only $c^{2}\cdot\varepsilon^{2}n^{2}$ bad edges, which, recalling that we started with $\frac{1}{\varepsilon}$ bad edges, matches the bound $\Omega(\varepsilon^{3}n^{2})$ on the price of uncertainty in Balcan et al. (2011). It turns out we can produce many more bad edges in our construction by doing the following. Note that the last layer in our construction is blue, and all layers between the frozen layers defined above and the last layer are now red. We now turn each vertex in the layer after the frozen layers blue, which is a valid uncertain best response by Proposition 3. We can now turn each successive layer blue in sequence until the last three layers are blue, red, and blue, respectively. By repeatedly switching the colors of the vertices in the layer after the frozen layers and propagating uncertain best responses, we can reach a state in which the boosting operation layers alternate between blue and red. Every edge between boosting operation layers is now bad, which makes
$\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{1-\left(\frac{1}{1+\frac{\varepsilon}{2}}\right)^{2}}\right)=\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{\frac{\varepsilon(4+\varepsilon)}{(2+\varepsilon)^{2}}}\right)=\Theta(\varepsilon n^{2})$
bad edges at the ending state. Recalling that we began with $\frac{1}{\varepsilon}$ bad edges, we have established that the price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ in the domain $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
@@ 4 Upper Bound
In addition to focusing only on the steps at which each vertex plays its first uncertain best response, we will also carefully construct and analyze a construction that restricts the degrees of new vertices and allows us to bound the number of edges in the graph.
@@ 4.1 Proof of Improved Upper Bound
Let us number the vertices in $P$ in order of their first best responses from first to last as $p_{1},p_{2},\dots,p_{m}$ for $m=|P|$. Then, for all integers $1\leq k\leq m$, define $S_{k}=S_{0}\cup\{p_{1},p_{2},\dots,p_{k}\}$. Notice that the first vertex $p_{1}$ can only change its color after several vertices in $S_{0}$ change colors, because $p_{1}$ must begin with only good edges. Now that we have set up our notation, we will prove some preliminary statements regarding the behavior of the consensus game.
@@
Clearly, $m=|P|<|V|=n$, so $m=O(n)$ and the upper bound holds. We will now prove the lower bound.
By construction, $V$ consists of the two disjoint subsets $S_{m}$ and $Q$. By Proposition 6, any vertices in $Q$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge. This means that we can delete these vertices along with all of their incident edges, which will never decrease the factor by which the number of bad edges in $G$ can be increased. We will take this a step further by deleting all vertices in $Q$ that do not share an edge with a vertex in $P$. By similar logic, this cannot decrease the PoU because all vertices in $V\setminus Q$ can still have at most as many good edges as before. Thus, the same sequence of uncertain best responses can be played to produce the same number of final bad edges even though we now have fewer vertices in $V$.
After all of our deletions, we are now left with a new set of vertices $Q^{\prime}\subset Q$ where every edge incident to a vertex in $Q^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=P$. Then, there exists a vertex $v\in P$ with the most edges connecting it to a vertex in $Q^{\prime}$. Let the number of such edges be $w$. We then construct a set $R$ of vertices $\{r_{1},r_{2},\dots,r_{w}\}$, and we can add new edges such that each vertex in $P$ connected by $j$ edges to vertices in $Q^{\prime}$ is connected to $r_{1},r_{2},\dots,r_{j}$. Finally, we delete all vertices in $Q^{\prime}$ and all of their incident edges. The factor by which the number bad edges in $G$ can be increased doesn’t change because all vertices that change colors still have the same number of good edges. Furthermore, $|R|=w\leq|Q^{\prime}|$, so the PoU cannot decrease. Finally, since $v$ always has at least $w=|R|$ good edges, it must have at least $|R|(1+\varepsilon)^{-1}=\Theta(|R|)$ bad edges when it plays its first uncertain best response. However, by construction, $v$ cannot have more than $|S_{m}|$ bad edges, so $|R|=O(|S_{m}|)$. Since the entire graph now only consists of vertices in $R$ and $S_{m}$, this shows that $n=\Theta(|S_{m}|)$.
Next, recall that $|S_{m}|=|S_{0}|+m$. By 5, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin with a bad neighbor is at most twice this quantity. So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well, and we thus have
$\Theta(n)$ $=|S_{m}|=|S_{0}|+m=O\left(\varepsilon^{-2}\right)+m.$
@@
$\Theta(n)$ $=o\left(n\right)+m\implies\Theta(n)=m.$
@@
Suppose that $A_{0}=\{3,2,2\}$, $\varepsilon=1/3$, and we wish to play a move on $z=3$. Then, $\lfloor(1+\varepsilon)z\rfloor=4$, and the possible multisets for $A_{1}$ are:
$\{3,3,1,1\},\quad\{3,2,1,1,1\},\;\;\text{and}\;\;\{2,2,1,1,1,1\}.$
@@
  • Pad $A_{k}$ with an sufficiently large (at least $\lfloor z(1+\varepsilon)\rfloor$) number of zeroes.
@@
  • Remove all remaining zeroes from $A_{k+1}$.
@@
Let $T_{k}$ denote the number of edges between two vertices in $S_{k}$. For every regular move that we play on $A_{k-1}$ with an integer $z\in A_{k-1}$, the sum $\sum A_{k}$ grows from $\sum A_{k-1}$ by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges between two vertices in $S_{k}$ grows from $S_{k-1}$ by $z$. This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$. Thus, we have
$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\sum A_{k}-\sum A_{k-1},$
as well as
$\varepsilon\left(T_{k}-T_{k-1}\right)-1=\varepsilon z-1<\lfloor\varepsilon z\rfloor,$
which by substitution implies the inequality
$\varepsilon\left(T_{k}-T_{k-1}\right)-1<\sum A_{k}-\sum A_{k-1}.$
@@
$\varepsilon\left(T_{m}-T_{0}\right)-m$ $\leq\sum A_{m}-\sum A_{0}.$
@@
$\varepsilon\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}.$
@@
$\varepsilon\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum A_{0}$
@@
By Proposition 11, $\varepsilon^{-1}\sum A_{0}$ must be $O(\varepsilon^{-1}n|S_{0}|)$. Since we assumed that $\varepsilon=\omega(n^{-1/3})$, this is dominated by the $O(\varepsilon^{2}n^{2}|S_{0}|)$ term. Also, since we must have $\Omega(\varepsilon^{-1})$ starting bad edges, $S_{0}$ is $\Omega(\varepsilon^{-1})$ as well. This means that $O(\varepsilon n^{2})$ must also be $O(\varepsilon^{2}n^{2}|S_{0}|)$, and then $\varepsilon=\omega(n^{-1/3})$ implies that $O(\varepsilon n^{2})$ will dominate $\Theta(\varepsilon^{-1}n)$. We are left with the following terms on the RHS:
@@
We will now show that $T_{0}$ cannot grow faster than $O(\varepsilon^{2}n^{2}|S_{0}|)$. By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$. By 5, we start with $O(\varepsilon^{-2})$ bad edges, so $T_{0}=O(\varepsilon^{-2})$. This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega(n^{-1/3})$. Then, $O(\varepsilon n^{2})$ is $O(\varepsilon^{2}n^{2}|S_{0}|)$ because $|S_{0}|=\Omega(\varepsilon^{-1})$. Our equation finally becomes
@@
Intuitively, playing a regular move on $\sum A_{k-1}$ with a large $z$ would result in a big increase in from $\sum A_{k-1}$ to $\sum A_{k}$. However, playing such a move might also create many small elements in $A_{k}$. Having the sum $\sum A_{k}$ spread out across many small elements will make increasing $\sum A_{k}$ in future moves more difficult. Recall that $\sum A_{k}^{2}$ denotes the sum of squares of the not necessarily distinct elements of $A_{k}$. We consider $\sum A_{k}^{2}$ as a way to measure how ”spread out” the multiset $A_{k}$ is, and this particular quantity is preferred over other measures because the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$ can be related to $\sum A_{k-1}$.
@@
Eq. 1 is true because the elements of $A_{m}$ must be nonnegative integers, and Eq. 2 follows from expanding both sides. We will only consider these inequalities (and not those with $m$ and 0 swapped) because we only need to prevent too many regular moves from being played with large $z$, which will naturally cause $\sum A_{k}^{2}$ to become small as $k$ increases.
We will begin by investigating the behavior of the sum of squares. Let the current multiset be $A_{k-1}$. Then, pick some element $z\in A_{k-1}$, and define $\zeta=\lfloor z(1+\varepsilon)\rfloor$. We can play a regular move on $A_{k-1}$ with $z$, which will involve incrementing the $\zeta$ elements $a_{1},a_{2},\dots,a_{\zeta}\in A_{k-1}$ to $a_{1}+1,a_{2}+1,\dots,a_{\zeta}+1\in A_{k}$. It is easy to see that the effect of such a move on the increase from $\sum A_{k-1}$ to $\sum A_{k}$ is
@@
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $=-z^{2}+\sum_{i=1}^{\zeta}\left((a_{i}+1)^{2}-a_{i}^{2}\right)$
@@
Eq. 3 and Eq. 4 together describe a single regular move. However, we can establish bounds upon these equations that are easier to analyze. Since $\varepsilon<1$, we have $\zeta=\lfloor z(1+\varepsilon)\rfloor\leq 2z$, so we can rewrite Eq. 4 as
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum_{i=1}^{\zeta}a_{i}-z^{2}+2z.$
Next, since $a_{1},a_{2},\dots,a_{\alpha}$ are all elements of $A_{k-1}$, their sum must be bounded by $\sum A_{k-1}$. By plugging this into our expression for the change from $\sum A_{k-1}^{2}$ to $\sum A_{k}^{2}$, we get
$\sum A_{k}^{2}-\sum A_{k-1}^{2}$ $\leq 2\sum A_{k-1}-z^{2}+2z.$ (5)
@@
Notice that the changes described in the right hand sides of Eq. 3, Eq. 4, and Eq. 5 are all independent of the current value of $\sum A_{k-1}^{2}$. Furthermore, the only constraint on the final sum of squares $\sum A_{m}^{2}$ is a lower bound provided by Eq. 1. This means that always taking our upper bound on $\sum A_{k}$ from Eq. 5 will never decrease $\sum A_{m}$. ∎
Now equipped with Proposition 13, we actually no longer need that $z\in A_{k-1}$. In fact, we no longer even need to consider $A_{k-1}$ as a multiset anymore. For the rest of the proof, we will simply assume that $z$ is a positive real number. The next step is to find an upper bound on the change from $\sum A_{k-1}$ to $\sum A_{k}$, which is
@@
$\varepsilon z^{\prime}$ $=\lfloor z(1+\varepsilon)\rfloor-z\leq z(1+\varepsilon)-z=\varepsilon z.$
Thus, $z^{\prime}\leq z$. By Eq. 5, smaller values of $z$ yield larger values of $\sum A_{k}^{2}$. This is exactly what we wanted. ∎
@@
$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{i}\right)-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}.$
We can then rewrite the $\sum A_{j}$ by plugging in Eq. 1 and Eq. 7, and we thus have
$\sum A_{m}^{2}$ $=\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\sum A_{0}+i+\varepsilon\sum_{j=1}^{i}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}$
$0$ $\leq\sum A_{0}^{2}+2\sum_{i=1}^{m}\left(\varepsilon\sum_{j=1}^{m}z_{j}\right)-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}$
$0$ $\leq\sum A_{0}^{2}+2\varepsilon m\sum_{i=1}^{m}z_{i}-\sum_{i=1}^{m}z_{i}^{2}+2\sum_{i=1}^{m}z_{i}.$ (8)
@@
Plugging this result into Eq. 8 yields the following quadratic polynomial in terms of $\sum_{i=1}^{m}z_{i}$ in the RHS:
$0$ $\leq\sum A_{0}^{2}+(2\varepsilon m+2)\sum_{i=1}^{m}z_{i}-m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}$
$0$ $\geq m^{-1}\left(\sum_{j=1}^{m}z_{j}\right)^{2}-(2\varepsilon m+2)\sum_{i=1}^{m}z_{i}-\sum A_{0}^{2}.$
Since the leading coefficient is positive and the quadratic polynomial is on the lesser side of the inequality, $\sum_{j=1}^{m}z_{j}$ must be less than the larger root of this polynomial. This larger root is obtained via the quadratic formula:
$\sum_{j=1}^{m}z_{j}\leq\frac{(2\varepsilon m+2)+\sqrt{(2\varepsilon m+2)^{2}+4m^{-1}\sum A_{0}^{2}}}{2m^{-1}}.$ (9)
@@
Suppose that $\varepsilon=\Omega(n^{-1/4})$. Then, $\sum_{i=1}^{m}z_{j}$ must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$.
@@
Since $\varepsilon=\Omega(n^{-1/4})=\Omega(m^{-1/4})$ by Lemma 7, the $2\varepsilon m$ in Eq. 9 must be $\omega(1)$. Thus, if we substitute asymptotic behavior into the RHS of Eq. 9, we get